kranti's mother is 4 times as old as kranti, after 5 years her mother will be 3 times as old as she will be then find their ages.
Answer by mananth(16946)

kranti's age =x
mother's age =4x
4x+5 = 3(x+5)
4x+5 =3x+15
4x-3x=15-5
x=10
Kranti's age = 10 years
mother's age = 4x = 40 years